Let P = [ -1 4 5] . Determine P^-1 using


[ 0 2 -3]


[ 0 0 8]


Cayley- Hamilton theorem. Further use P^-1 to express (x1, x2, x3) in terms of (-1,0,0), (4,2,0), ( 5,-3,8)

1
Expert's answer
2021-12-02T12:17:02-0500

"P=\begin{bmatrix}
-1&4&5
\\
0&2&-3
\\
0&0&8
\end{bmatrix}"

"p(\lambda)=\det (P-\lambda I)=\begin{vmatrix}
-1-\lambda&4&5
\\ 0&2-\lambda &-3
\\
0&0&8-\lambda 
\end{vmatrix}
=(-1-\lambda )(2-\lambda )(8-\lambda)"

"p(\lambda )=-x^3+9x^2-6x-16"


"-P^3+9P^2-6P-16I=0"

"P(-P^2+9P-6I)=16I"

"16P^{-1}=-P^2+9P-6I=-\begin{bmatrix}
1&4&23
\\
0&4&-30
\\
0&0&64
\end{bmatrix}+9\begin{bmatrix}
-1&4&5
\\
0&2&-3
\\
0&0&8
\end{bmatrix}-\begin{bmatrix}
6&0&0
\\
0&6&0
\\
0&0&6
\end{bmatrix}"


"16P^{-1}=\begin{bmatrix}
-16&32&22
\\
0&8&3
\\
0&0&2
\end{bmatrix}"


"P^{-1}=\frac{1}{16}\begin{bmatrix}
-16&32&22
\\
0&8&3
\\
0&0&2
\end{bmatrix}"


Let "\begin{bmatrix}
x_1 \\ x_2\\ x_3
\end{bmatrix}=a\cdot \begin{bmatrix}-1\\0\\0 \end{bmatrix}+ b\cdot \begin{bmatrix}4\\2\\0 \end{bmatrix}+ c\cdot \begin{bmatrix}5\\-3\\8 \end{bmatrix}=P\begin{bmatrix}a\\b\\c\end{bmatrix}" .


Then "\begin{bmatrix}
a\\b\\c
\end{bmatrix}=P^{-1}\begin{bmatrix}x_1\\x_2\\x_3\end{bmatrix}=
\begin{bmatrix}
-x_1+2x_2+\tfrac{11}{8} x_3
\\ \tfrac{1}{2}x_2+\tfrac{3}{16}x_3
\\
\tfrac{1}{8}x_3
\end{bmatrix}" .


So, "\begin{bmatrix}
x_1 \\ x_2\\ x_3
\end{bmatrix}=(-x_1+2x_2+\tfrac{11}{8}x_3)\cdot \begin{bmatrix}-1\\0\\0 \end{bmatrix}+ (\tfrac{1}{2}x_2+\tfrac{3}{16}x_3)\cdot \begin{bmatrix}4\\2\\0 \end{bmatrix}+ \tfrac{1}{8}x_3\cdot \begin{bmatrix}5\\-3\\8 \end{bmatrix}" .




Answers:

"P^{-1}=\frac{1}{16}\begin{bmatrix}
-16&32&22
\\
0&8&3
\\
0&0&2
\end{bmatrix}"


"\begin{bmatrix}
x_1 \\ x_2\\ x_3
\end{bmatrix}=(-x_1+2x_2+\tfrac{11}{8}x_3)\cdot \begin{bmatrix}-1\\0\\0 \end{bmatrix}+ (\tfrac{1}{2}x_2+\tfrac{3}{16}x_3)\cdot \begin{bmatrix}4\\2\\0 \end{bmatrix}+ \tfrac{1}{8}x_3\cdot \begin{bmatrix}5\\-3\\8 \end{bmatrix}"